# Aligning with Logic: Measuring, Evaluating and Improving Logical Consistency in Large Language Models

### v2

@@ 3.2 Results and Analysis
Performance of Different (Families of) Models. As shown in Table 1, recent LLMs like Gemma2-9B and Phi-3-medium demonstrate stronger overall consistency compared to earlier models. In particular, models such as Deepseek-chat, Phi-3-medium, and Gemma-2-9B perform well across all three evaluated consistency dimensions. However, it is important to note that strong performance in one aspect of consistency does not guarantee similar performance in others. For example, while Mistral-7B excels in transitivity, its performance in other consistency aspects is weaker.
The Phi-3 family consistently shows high logical consistency, which we hypothesize is due to its reliance on synthesized data which is cleaner and less self-contradictory. This type of data likely reduces internal conflicts, resulting in more consistent logical behavior. However, no strong correlation is observed between logical consistency and human agreement accuracy. For instance, while Gemma-2-9B and Llama-3-8B achieve similar levels of human agreement, Gemma-2-9B shows higher consistency in its outputs. This is further confirmed by the low correlations (below $0.05$) between transitivity consistency and human preference accuracy across all datasets.
@@
Impact of CoT Prompting. We also investigated the effect of CoT prompting on logical consistency. Surprisingly, CoT reasoning did not generally improve consistency, and in some cases, it led to a decrease in transitivity performance. We hypothesize that this decline may be due to the additional reasoning tokens generated during CoT prompting, which could introduce variations in the judgement standards used for different comparisons. When a model’s understanding of the preferential relations is not uniform, it may produce inconsistent (non-transitive) outcomes. This suggests that CoT prompting, while beneficial for complex reasoning, might introduce unintended inconsistencies in certain logical judgement tasks.
@@ 3.3 Consistency and Reliability
Transitivity as a Proxy for Self-Agreement. We next investigated the relationship between transitivity and self-agreement across different LLMs. As shown in Figure 4, there are strong correlations between transitivity and self-agreement across all three datasets, regardless of the task’s level of subjectiveness. Self-agreement indicates the model’s internal consistency and robustness, as a reliable model should not fluctuate significantly in its responses. Higher self-agreement suggests that the model exhibits a stable understanding of the underlying relations in the input. Given that transitivity captures this self-consistency, we argue that transitivity serves as a useful proxy for evaluating the global reliability of LLMs.
Commutativity Correlates Strongly with Human Preferences. For each task, we paraphrased the comparison prompt template 10 times using GPT-4-turbo to explore the sensitivity of LLMs to input variations. Despite these paraphrases being semantically equivalent, they produced different performance outcomes, as shown in Figure 5.
We found strong correlations between commutativity and human agreement rates across nearly all datasets and models. This finding is not surprising, as a lack of commutativity often indicates a strong positional bias. These results align with previous research (Zhou et al., 2024), which demonstrated that positional bias can significantly impact alignment with human judgments.
@@
Figure 5: Commutativity shows a generally strong correlation with human preference agreement rates across various LLMs and all three datasets.
4 A Data Refinement and Augmentation Approach to Improve Logical Consistency in LLMs
@@
Our analysis so far has demonstrated that many LLMs exhibit varying degrees of logical (in)consistency when making judgements. To address this challenge, we propose a data refinement and augmentation framework that aims to mitigate this phenomenon. The idea is to filter out self-conflicting information from noisy preference data and generate additional conflict-free pairwise comparisons. This approach aims to enhance the logical consistency of LLMs while maintaining alignment with human preferences, enabling them to function more reliably as logical operators.
Motivation. Previous work has proven the benefits of using pairwise comparisons derived from preference rankings to train LLMs for a better understanding of preference orderings (Song et al., 2024). Additionally, Asai & Hajishirzi (2020) showed that incorporating logically derived counterparts of pairwise annotations improves consistency in knowledge-intensive QA tasks. However, real-world preference data are often noisy and self-contradictory (Chowdhery et al., 2022; Ouyang et al., 2022a), especially in more subjective tasks (Bai et al., 2022). The inter-rater agreement rate typically falls between $60\%$ to $80\%$ for preference or evaluation datasets. Prior work (Wang et al., 2024a) has shown that these self-conflicting annotations reduce the efficiency of learning preferences, and we hypothesize that such annotations also contribute to the logical inconsistencies in trained models.
Inspired by the RLHF process, where a reward model is used to establish a complete order of responses, we propose first estimating a coherent ranking from the noisy data. This refined ranking can then be augmented with additional conflict-free pairwise comparisons to align LLMs more effectively. This method is expected to be more efficient than aligning models to noisy and inconsistent annotations directly.
Table 2: Data statistics at each augmentation stages. # Data refers to data size and AvgComp/Inst for average comparisons per instance. Aug. data represents Augmented data, and Neg. denotes further augmentation with negated relation comparisons.
@@
Aug. data 1 1 30.9K 13.2
Aug. data+Neg. 1 1 61.8K 26.4
Table 3: Instruction-tuning (IT) of Llama-3 Instruct (8B) using different dataset variants (see Section 4.2). Only training on negated relations can enhance negation invariance.
@@ H. $s_{tran}(5)$ $s_{comm}$ $s_{neg}$
IT Perturbed data 70.3 91.9 88.4 61.0
IT Augmented data 70.1 95.4 91.2 60.8
IT Augmented data + Neg. 69.7 95.2 90.5 87.9
4.1 Estimating Rankings from Noisy Pairwise Annotations Using Win-Loss Rate
Estimating global rankings from noisy pairwise annotations is essentially a rank aggregation problem. We use the win-loss rate method due to its simplicity and two key advantages: 1) it performs well with partial and sparse comparisons, which is common in real-world preference datasets, and 2) it remains unaffected by the order in which comparisons are presented. While other rank aggregation methods like ELO or the Bradley-Terry model (Bradley & Terry, 1952) could be explored for improved results, our choice serves as a proof-of-concept rather than an optimization for best performance.
The win-loss rate for each item is calculated as the number of comparisons it wins minus the number of comparisons it loses and then divided by the number of comparisons it participates in. Following that, as shown in Figure 6, we rank the item by the value of its win-loss rate. This way, we can aggregate a full or partial ranking. We then split the ranking into a self-consistent pairwise comparison set. We can further augment the dataset by adding the comparisons using the negated relation.
@@ 4.2 Experiments
Experimental Setup. We use the Summarize From Feedback dataset (Mostafazadeh et al., 2016b), where human annotators made pairwise comparisons between two summaries based on qualitative factors. More details are available in Appendix §B. Given that the dataset annotations are sparse and relatively clean, we simulate noise by randomly flipping $10\%$ of the training labels. Our data refinement and augmentation techniques are then applied to improve the consistency and quantity of pairwise comparisons, as shown in Table 3.
To assess the effectiveness of our approach, we instruction-tuned three Meta-Llama-3-8B-Instruct models on: 1) the flipped/perturbed data, 2) the refined and augmented dataset, and 3) the ‘further augmented’ dataset with additional negated relation comparisons. Training parameters are detailed in Appendix § E. We randomly sample 200 instances from the test set and evaluate all models on this subset. We assess the impact of the data augmentation on logical consistency and performance.
Results and Findings. The results shown in Table 3 highlight three key findings: 1) Zero-shot inference shows considerable logical inconsistency. However, training on perturbed data can already substantially improve both human preference alignment and logical consistency. 2) Training with the refined and augmented dataset significantly improves transitivity and commutativity, while maintaining strong human alignment. 3) Only training further on negated relations improves the model performance in negation invariance. However, adding negated relations to the broader dataset may introduce distractions and cause a forgetting effect (Luo et al., 2023), resulting in a slight reduction in performance on other logical properties. Overall, the findings support the effectiveness of our data refinement and augmentation framework in improving the logical consistency of LLMs.
5 Impact of Logical Consistency on Downstream Applications
LLMs are increasingly used as logical operators in high-level algorithms due to their ability to process and understand text semantics. For instance, Qin et al. (2024) use LLMs to enhance document reranking in information retrieval systems. Similarly, Guo et al. (2024) and Yang et al. (2024) utilize LLMs as optimizers in prompt-tuning algorithms, while Liu et al. (2024) and Liusie et al. (2024b) employ LLMs as pairwise comparators to aggregate preference rankings. When LLMs are used as logical operators, maintaining a high level of logical consistency is critical to ensure predictable and efficient decision-making. In this section, we examine how logical consistency influences the performance of LLM-based algorithms in such ‘logically grounded’ tasks.
@@ H. $s_{tran}(5)$ $s_{comm}$ PairS PairS calibrated
5.1 Application Example: Preference Ranking Algorithm
We illustrate the impact of logical consistency through the Pairwise-Preference Search (PairS) algorithm proposed by Liu et al. (2024). PairS is a sorting-based rank aggregation method that uses LLMs as pairwise evaluators (i.e., it is a particular ‘LLM-as-a-judge’ algorithm), comparing items based on specific attributes using a merge-sort approach. Its performance is measured by comparing the LLM-generated rankings with human judgments via Spearman’s correlation. Sorting algorithms depend heavily on logical properties such as transitivity and commutativity. PairS assumes that LLMs used as evaluators have near-perfect transitivity and commutativity for optimal ranking results.
To evaluate this, we conduct controlled experiments on the coherence aspect of SummEval, using various LLMs with similar accuracy to human judgments. The results are shown in Table 4. As LLMs often exhibit positional bias, we follow Liu et al. (2024) by reporting both raw and calibrated results. The calibration, performed as described in Zheng et al. (2023), averages the evaluation probabilities across both possible pairwise orders. While this technique increases inference cost, it reduces bias by balancing positional preferences.
Our findings from Table 4 reveal two key insights: 1) Although Phi-3-mini has slightly lower accuracy with human judgments than GPT-3.5-turbo, its stronger transitivity leads to better ranking performance with or without calibration. 2) There is a clear correlation between an LLM’s commutativity and the performance gains from calibration in the PairS algorithm. LLMs that are more commutative rely less on calibration to achieve optimal performance, and therefore, require less computations.
@@ 6 Related Work
Despite the wealth of knowledge regarding the logical consistency of LLMs within specific domains, most studies have concentrated on first-order relations between only two or three statements. Consequently, there is a notable gap in research addressing the consistency and reliability of LLMs when applied to more complex decision-making scenarios or the evaluation of multiple items simultaneously. This limitation suggests a pressing need for further investigation into how LLMs can maintain coherence and consistency in broader, more comprehensive contexts, which is essential for their deployment in practical applications.
@@ 7 Conclusion
In this work, we investigated the critical role of logical consistency in enhancing the reliability and trustworthiness of LLMs, especially in their roles as decision-makers and reasoners. We introduced a general framework for quantifying three key properties of logical consistency: transitivity, commutativity, and negation invariance. Through comprehensive evaluation, we demonstrated that these consistency measures serve as strong proxies for assessing the overall reliability of LLMs.
To improve the logical consistency of LLMs, we then proposed a data refinement and augmentation framework, which effectively reduces logical inconsistencies at the source: directly in the data. Our experimental results showed that models trained on this refined and augmented data achieve improved logical consistency without compromising alignment with human preferences. Furthermore, we demonstrated that integrating LLMs with higher logical consistency into a logic-driven algorithm improves both the algorithm’s global performance and computational efficiency. This highlights the practical benefits of ensuring logical coherence in LLMs when applied to downstream tasks.
In summary, our work emphasizes the importance of logical consistency as a complementary factor to human alignment in the development of more reliable LLMs. We hope this research will encourage the broader community to recognize the importance of logical consistency as a fundamental aspect of any work aiming towards improved trustworthiness and reliability.
@@ Appendix A Transitivity calculation
2:     $visited\leftarrow visited\cup\{v\}$, $recStack\leftarrow recStack\cup\{v\}$
3:     for each $u$ in $G[v]$ do
4:         if $u\notin visited$ and CycleUtil($u$, visited, recStack) then
5:              Return True
6:         else if $u\in recStack$ then
7:              Return True
8:         end if
9:     end for
10:     $recStack\leftarrow recStack\setminus\{v\}$
11:     Return False
12:end function
@@ Appendix C Large Language Models in Evaluation
API-Based LLMs. The GPT-3.5-turbo-0125 model is part of OpenAI’s GPT-3.5 series, which is renowned for its strong performance in natural language understanding and generation. Deepseek-chat (DeepSeek-AI, 2024) is another API-based LLM developed by DeepSeek.
@@ Appendix D Additional Related Work
Improving Consistency of LLMs. Asai & Hajishirzi (2020) proposed an augmentation method and a training regularization term to improve the logical consistency of language models for QA tasks. Minervini & Riedel (2018) proposed an adversarial training procedure to improve the robustness of NLI models to adversarial examples. Wang & Henao (2021) focus on improving the paraphrasing consistency during training for the task of low-resource Named Entity Recognition (NER).
@@
Appendix E Instruction-Tuning: Training Details

### v3

@@ 3.2 Results and Analysis
Performance of Different (Families of) Models. As shown in Table 1, recent LLMs like Gemma2-9B and Phi-3-medium demonstrate stronger overall consistency compared to earlier models. In particular, models such as Deepseek-chat-v2, Phi-3-medium, and Gemma-2-9B perform well across all three evaluated consistency dimensions. However, it is important to note that strong performance in one aspect of consistency does not guarantee similar performance in others. For example, while Mistral-7B excels in transitivity, its performance in other consistency aspects is weaker. The Phi-3 family consistently shows high logical consistency, which we hypothesize is due to its reliance on synthesized data which is cleaner and less self-contradictory.
@@
Impact of CoT Prompting. We also investigated the effect of Chain-of-Thought (CoT) prompting (Wei et al., 2022) on logical preference consistency. Surprisingly, CoT reasoning did not generally improve consistency, and in some cases, it led to a decrease in transitivity performance. We hypothesize that this decline may be due to the additional CoT reasoning tokens could introduce variations in the judgment standards used for different comparisons. When the understanding of the preferential relations is not uniform, it may produce inconsistent (non-transitive) outcomes. This suggests that CoT prompting, while beneficial for complex reasoning, might introduce unintended inconsistencies in certain logical judgment tasks.
Figure 5: Commutativity shows a generally strong correlation with human preference across various LLMs and datasets.
@@ 3.3 Consistency and Reliability
Transitivity as a Proxy for Self-Agreement. As shown in Figure 4, there are strong correlations between transitivity and self-agreement across all three datasets, regardless of the task’s level of subjectiveness. Self-agreement indicates the model’s internal consistency and robustness, as a reliable model should not fluctuate significantly in its responses. Higher self-agreement suggests that the model exhibits a stable understanding of the underlying relations in the input. Given that transitivity captures this self-consistency, we argue that transitivity serves as a useful proxy for evaluating the local robustness of LLMs.
Commutativity Correlates Strongly with Human Preferences. For each task, we paraphrased the comparison prompt template 10 times using GPT-4-turbo to explore the sensitivity of LLMs to input variations. Despite these paraphrases being semantically equivalent, they produced different performance outcomes, as shown in Figure 5. We observe strong correlations between commutativity and human agreement rates across nearly all datasets and models. This finding is not surprising, as a lack of commutativity often indicates a strong positional bias. These results align with previous research (Zhou et al., 2024), which demonstrated that positional bias can significantly impact alignment with human judgments.
@@
4 Improve Logical Preference Consistency in LLMs via REPAIR
@@
Our analysis so far has demonstrated that LLMs exhibit inconsistent logical behavior in making preferential judgments. To address this, we propose REPAIR (Ranking Estimation and Preference Augmentation through Information Refinement), a framework that aims to mitigate these inconsistencies. REPAIR first estimates a ranking from noisy preference data, then generates additional conflict-free pairwise comparisons. This approach enhances logical preference coherence while maintaining alignment with human preferences, improving LLMs’ reliability as logical operators. See Appendix §K for detailed motivation.
Table 2: Statistics at each augmentation stage. # Data means data size and AvgComp/Inst for average comparisons per instance. Aug. data represents Augmented data and Neg. denotes further augmentation with negated relation comparisons.
@@
REPAIR-ed data 1 1 30.9K 13.2
REPAIR-ed data+Neg. 1 1 61.8K 26.4
Table 3: Instruction-tuning (IT) of Llama-3 Instruct (8B) using different variants (see Section 4.2). Only training on negated relations can enhance negation invariance.
@@ H. $s_{tran}(5)$ $s_{comm}$ $s_{neg}$
Perturbed data 70.3 91.9 88.4 61.0
REPAIR-ed data 70.1 95.4 91.2 60.8
REPAIR-ed data + Neg. 69.7 95.2 90.5 87.9
4.1 Estimating Rankings from Noisy Pairwise Data
Estimating global rankings from noisy pairwise annotations is essentially a rank aggregation problem. We use the win-loss rate method due to its simplicity and two key advantages: 1) it performs well with partial and sparse comparisons, which is common in real-world preference datasets, and 2) it remains unaffected by the order in which comparisons are presented. In Appendix §N, we compare the performance of alternative rank aggregation methods, such as the ELO rating and the Bradley-Terry model, within the REPAIR framework. Additionally, we provide a detailed analysis of their respective strengths and limitations.
The win-loss rate for each item is calculated as the number of comparisons it wins minus the number of comparisons it loses and then divided by the number of comparisons it participates in. Following that, as shown in Figure 6, we rank the item by the value of its win-loss rate. Through this way, a full or partial ranking is aggregated. We then extrapolate the ranking into a self-consistent pairwise comparison set, which can be further augmented by adding comparisons with the negated relation.
@@ 4.2 Experiments
Experimental Setup. We use the Summarize From Feedback dataset (Mostafazadeh et al., 2016b), where annotators made pairwise comparisons between two summaries based on qualitative factors. Further details about this dataset can be found in Appendix §B. Additionally, we provide experiments on another dataset in Appendix §L. Given that annotations are sparse and relatively clean, we simulate noise by randomly flipping $10\%$ of the training labels. Our data refinement and augmentation techniques are then applied to improve the consistency and quantity of pairwise comparisons, as shown in Table 3.
To assess the effectiveness of REPAIR, we instruction-tuned three Meta-Llama-3-8B-Instruct models on: 1) the flipped/perturbed data, 2) the refined and augmented dataset, denoted as REPAIR-ed data, and 3) The REPAIR-ed dataset with additional negated relation comparisons. Training hyperparameters are detailed in Appendix § F. For evaluation, we randomly sample 200 instances from the test set and evaluate all models on this subset. The impact of the data augmentation is assessed in terms of logical consistency and model performance.
Results and Findings. Results in Table 3 highlight three key findings: 1) Zero-shot inference shows considerable logical inconsistency. However, training on perturbed data can already substantially improve both human preference alignment and consistency. 2) Training with the REPAIR-ed dataset significantly improves transitivity and commutativity, while maintaining strong human alignment. 3) Only training further on negated relations improves the model performance in negation invariance. However, adding negated relations to the broader dataset may introduce distractions and cause a forgetting effect (Luo et al., 2023), resulting in a slight reduction in performance on other logical properties. Overall, the findings support the effectiveness of REPAIR in improving the logical preference consistency of LLMs.
5 Impact of Logical Preference Consistency on Downstream Applications
LLMs are increasingly used as logical operators in high-level algorithms due to their ability to process and understand text semantics. For instance, Qin et al. (2024) use LLMs to enhance document reranking in information retrieval systems. Similarly, Guo et al. (2024) and Yang et al. (2024) utilize LLMs as optimizers in prompt-tuning algorithms, while Liu et al. (2024) and Liusie et al. (2024b) employ LLMs as pairwise comparators to aggregate preference rankings. When LLMs are used as logical operators, maintaining a high level of logical consistency is critical to ensure predictable and efficient decision-making. In this section, we examine how logical preference consistency influences the performance of LLM-based algorithms in such ‘logically grounded’ tasks.
@@ H. $s_{tran}(5)$ $s_{comm}$ PairS PairS calibrated
We illustrate the impact of logical preference consistency through the Pairwise-Preference Search (PairS) algorithm proposed by Liu et al. (2024). PairS is a sorting-based rank aggregation method that uses LLMs as pairwise evaluators (i.e., a particular ‘LLM-as-a-judge’ algorithm), comparing items based on specific attributes using a merge-sort approach. Its performance is measured by comparing the LLM-generated rankings with human judgments via Spearman’s correlation. Sorting algorithms depend heavily on logical properties. PairS assumes that LLMs used as comparators have near-perfect transitivity and commutativity for optimal ranking results.
To evaluate this, we conduct controlled experiments on the coherence aspect of SummEval, using various LLMs with similar accuracy to human judgments. As shown in Table 4, we follow Liu et al. (2024) by reporting both raw and calibrated results. The calibration, performed as described in Zheng et al. (2023), averages the evaluation probabilities across both possible pairwise orders. While this technique increases inference cost, it reduces bias by balancing positional preferences. Our findings from Table 4 reveal two key insights: 1) Although Phi-3-mini has slightly lower accuracy with human judgments than GPT-3.5-turbo, its stronger transitivity leads to better ranking performance with or without calibration. 2) There is a clear correlation between an LLM’s commutativity and the performance gains from calibration in the PairS algorithm. LLMs that are more commutative rely less on calibration to achieve optimal performance, and therefore, require less computations.
@@ 6 Related Work
Despite the wealth of knowledge regarding the logical consistency of LLMs within specific domains, most studies have concentrated on first-order relations—logical connections or implications directly linking two or three individual statements. Consequently, there is a notable gap in research addressing the consistency and reliability of LLMs when applied to more complex decision-making scenarios or the evaluation of multiple items simultaneously. This limitation suggests a pressing need for further investigation into how LLMs can maintain coherence and consistency in broader, more comprehensive contexts, which is essential for their deployment in practical applications.
@@ 7 Conclusion
In this work, we explored the critical role of logical preference consistency in enhancing the reliability and trustworthiness of LLMs, introducing a framework to quantify three key properties: transitivity, commutativity, and negation invariance, which serve as strong proxies for assessing overall reliability. To address logical inconsistencies, we proposed REPAIR, a data refinement and augmentation framework, demonstrating that models trained on REPAIR-ed data achieve higher logical preference consistency without compromising human alignment. Additionally, integrating logically consistent LLMs into logic-driven algorithms enhanced global performance and computational efficiency, underscoring the practical benefits of logical coherence in downstream tasks. Our findings highlight logical consistency as a complementary factor to human alignment in developing more reliable LLMs, encouraging its recognition as a fundamental aspect of improving trustworthiness and reliability.
Impact Statement
There are several potential societal consequences of our work. By ensuring logical coherence in LLMs, we hope to contribute to the development of more reliable and trustworthy AI systems. This could have positive implications for various sectors, including healthcare, finance, and education, where decision-making and reasoning tasks are critical. However, it is important to consider the ethical aspects of deploying such systems. Ensuring logical consistency does not automatically guarantee fairness, transparency, or accountability. Therefore, we encourage the broader community to continue exploring these ethical dimensions in parallel with technical advancements. In summary, while our work aims to advance the field of artificial intelligence by improving the logical consistency of LLMs, we recognize the need for ongoing discussions and research to address the ethical and societal implications of deploying AI systems.
@@ Appendix A Transitivity Calculation
2:  $visited\leftarrow visited\cup\{v\}$
3:  $recStack\leftarrow recStack\cup\{v\}$
4:  for each $u$ in $G[v]$ do
5:    if $u\notin visited\land\text{CycleUtil}(u,visited,recStack)$ then
6:      Return True
7:    else if $u\in recStack$ then
8:      Return True
9:    end if
10:  end for
11:  $recStack\leftarrow recStack\setminus\{v\}$
12:  Return False
@@ Appendix B Dataset
MS MARCO. (Nguyen et al., 2016) is a widely used large-scale benchmark for training and evaluating retrieval-based question answering systems. Following the methodology of Sun et al. (2023), we leverage the dataset for a document reranking task. Specifically, they randomly sample a subset of queries from the MS MARCO training set, retrieving 10 candidate passages for each query using BM25. Ranking preferences for these passages were distilled using GPT-3.5-turbo. In this work, we randomly sample 1K examples from their query subset and corresponding ranking labels. We then split them into a training set of 800 examples and a test set of 200 examples.
@@ Appendix C Large Language Models in Evaluation
API-Based LLMs. The GPT-3.5-turbo-0125 model is part of OpenAI’s GPT-3.5 series, which is renowned for its strong performance in natural language understanding and generation. Deepseek-chat-v2 (DeepSeek-AI, 2024) is another API-based LLM developed by DeepSeek.
@@ Appendix D Additional Related Work
Improving Consistency of LLMs. Asai & Hajishirzi (2020) proposed an augmentation method and a training regularization term to improve the logical consistency of language models for QA tasks. Minervini & Riedel (2018) proposed an adversarial training procedure to improve the robustness of NLI models to adversarial examples. Wang & Henao (2021) focus on improving the paraphrasing consistency during training for the task of low-resource Named Entity Recognition (NER). Li et al. (2019) proposed a framework to regularize the loss function when training a neural model, based on logic rules to improve the logical consistency.
@@
Appendix E Distinguishing Logical Transitivity from Social Network Analysis (SNA) Transitivity
Transitivity is a well-studied concept in Social Network Analysis (SNA). However, the definition and purpose for transitivity in our framework diverge significantly from those in SNA. Below, we highlight the key distinctions:
Motivations. In SNA, transitivity measures the tendency of nodes in a network to form tightly connected groups. It is based on real-world social patterns, where ”friends of friends” are likely to be friends, leading to clustered communities. In contrast, our logical consistency framework employs transitivity to assess self-conflict in relation graphs generated by models or human judgments. Here, transitivity serves as an indicator of logical coherence rather than network clustering.
Definition. In graph theory, transitivity quantifies the formation of closed triangles (fully connected triplets), indicating the probability that two neighbors of a node are also connected. In our framework, however, it measures the extent to which relation graphs are acyclic. Specifically, it evaluates the probability that a randomly selected set of nodes forms an acyclic structure, extending beyond simple triplet closures.
Application and Scenarios. SNA transitivity is primarily applied to undirected social network graphs, while our logical transitivity operates exclusively on directed relation graphs. SNA transitivity focuses on the closure of triplets, whereas our logical transitivity evaluates logical circulations spanning 3 to K nodes, with K being the size of the sampled sub-graph. Extending SNA measures to higher-order relationships is nontrivial and computationally inefficient.
To illustrate these differences, we provide a detailed analysis of popular SNA transitivity measures:
  1. 1.
    Clustering Coefficient (Wasserman, 1994): Global clustering coefficient measures
    $C=\frac{\text{Number of Closed Triplets}}{\text{Number of All Triplets (closedand open)}},$
    and local clustering coefficient measures
    $C_{i}=\frac{\text{Number of closed triplets including node $i$}}{\text{All triplets including node $i$}}.$
    While applicable to directed graphs, clustering coefficient still measures the ratio between closed triplets and total triplets instead of circular relations and cannot be applied to larger structures.
  2. 2.
    Triad Census (Batagelj & Mrvar, 2001): Counts the relative frequency of 16 possible triadic configurations in directed graphs. However, this method is inherently constrained to triads.
  3. 3.
    E-I (External-Internal) Index (Krackhardt & Stern, 1988): Measures the ratio of external to internal ties within a given partition, capturing group closure rather than logical acyclicity.
Appendix F Instruction-Tuning: Training Details